# Auto-bidding in real-time auctions via Oracle Imitation Learning

### v2

@@ 1. Introduction
3. Experimental validation. Using a simulator based on the Alimama dataset (Xu et al., [n. d.]), we show that OIL-trained agents achieve near-oracle performance and surpass other approaches, including PPO, BC, IQL, and LP.
@@ 2. Related work
Auto-bidding in second price auctions. A core component of advertisement bidding is estimating the utility value $V=K\times\mu$ of each IO, where $K$ is the target CPA and $\mu$ is the acquisition rate (Xu et al., 2016). $K$ is a parameter selected by the advertiser, stating how much it is ready to pay for an acquisition, while $\mu$ can be a function of the advertiser, the potential customer and the context (Yan et al., 2009; Wu et al., 2009; Yuan et al., 2013; Zhang et al., 2014). It can be either estimated by the advertisement platform or by the advertiser itself (Diemert Eustache, Meynet Julien et al., 2017). $V$ defines the fair value of an IO (Xu et al., 2016). Theoretical results prove that bidding equal to the fair value (truth-telling) is optimal for second-price auctions, under the assumption of infinite budget (Krishna, 2009; Roughgarden, 2010). However, when the budget is limited, it might be convenient to bid below the fair value, to increase the acquisition efficiency (Gong et al., 2023), or even above the fair value, if the advertiser can accept exceeding the target CPA to increase the number of acquisitions. The problem is more complex in the presence of multiple advertisement slots for a single IO. In this case, the efficiency of these slots varies with the position, and techniques such as bid shading can be employed to maximize the number of acquisitions (Zhang et al., 2021; Gong et al., 2023). Previous work has focused on optimal advertisement bidding when the fair value of an IO is unknown (Ren et al., 2017; Gligorijevic et al., 2020; Bompaire et al., 2021). In this work, instead, we consider the framework defined by Xu et al. (Xu et al., [n. d.]), in which the advertisement platform provides an estimation of the acquisition probability, and the advertiser optimizes its bids accordingly, as a function of its budget, target CPA and its understanding of the advertisement traffic patterns. We chose this specific problem because it can be tackled using the recently released Alimama dataset, which, with its more than 500 million recorded bids (Xu et al., [n. d.]), is the largest open-source advertisement bidding dataset to this date.
@@ 3. Dataset and simulator
Figure 2. Example IOs from the Alimama-dense dataset. A. Conversion probabilities for a single advertiser and advertisement campaign. B. Conversion probability at one time step (red dotted line, A) as a function of their lowest price (3rd slot). IOs above the dotted line have an expected CPA lower than the target ($K=8$).
Open-source bidding datasets have historically been limited in size and feature diversity (Ou et al., 2023), and usually focus on predicting the click and conversion probability (Diemert Eustache, Meynet Julien et al., 2017). Consequently, no common benchmark for auto-bidding agents exists, and recent works have relied on private datasets and closed-source algorithm implementations (Wang et al., 2017a; Zhang et al., 2021; Zhou et al., 2021; He et al., 2021; Wang et al., 2022; Wen et al., 2022). This tendency complicates the comparison between methods (Ou et al., 2023). However, recently Alimama has released a large scale dataset with more than 500 million advertisement bids showcasing realistic advertisement traffic patterns (Xu et al., [n. d.]).
Dataset. The dataset comprises two datasets, Alimama-dense and Alimama-sparse, spanning 21 campaign periods of 48 time steps each. For each time step and IO, bids from 48 advertisers are recorded, competing for three advertisement slots assigned via second-price auction. This means that the advertisers presenting the three highest bids win the three slots, paying a price equal to the next highest bid (second to fourth). The main distinction between the two datasets is the average conversion probability, which is one order of magnitude lower in the sparse dataset, resulting in fewer conversions (hence the name ”sparse”). Each advertiser is characterized by a fixed total campaign budget and a fixed target CPA, and bids according to its specific strategy. The dataset also provides the expected conversion probability for each IO, that was made available to the advertisers by the platform at bidding time. Conversion probabilities differ across advertisers for the same IO, reflecting varying product appeal to different potential customers. Advertisers face stochastic advertisement traffic (Fig. 2 A). To avoid both overbidding (and missing out on valuable IOs later due to lack of budget) and underbidding (and ending up with unspent budget that they had reserved for valuable IOs which never materialized), the advertisers need to adjust their strategy to the advertisement traffic they expect in the future. The conversion probabilities of the IOs at the same time step and for the same advertiser present a fairly spread distribution (Fig. 2 B), with only a fraction being convenient for the advertiser (i.e., the bid necessary to win the IO leads to a lower CPA than the target).
@@
Simulator. To train auto-bidding agents, we developed an advertisement campaign simulator based on the Alimama dataset (Fig. 3). The process is as follows: 1) the auto-bidding agent impersonates one of the 48 advertiser, sampled at random, and is assigned a budget and a target CPA. 2) At each time step, the simulator provides estimated conversion probabilities for available IOs (typically 1000 10,000 per time step). 3) The auto-bidding agent places bids without observing those of the competitors (closed-envelope auction). 4) Competitor bids are extracted from the dataset, assuming that they are not influenced by the auto-bidding agent’s behavior. 5) Slot allocation and pricing are determined via second-price auction.
@@ 4.2. Heuristic algorithm
Exact solutions to Problem 2 are computationally infeasible for campaigns involving hundreds of thousands of IOs, such as the ones in the Alimama dataset. Instead, we employ a greedy heuristic approximation. Inspired by knapsack problem solutions, the algorithm ranks slots by efficiency (e.g., conversion probability over cost, $\mu_{t,i}/k_{t,i,d}$) and selects slots until the budget is exhausted. Expected conversion probabilities and costs are adjusted for exposure probabilities since conversions and costs only occur when ads are displayed (effective conversion probability $\tilde{\mu}_{t,i,d}=\mu_{t,i}h_{d}$ and effective cost $\tilde{k}_{t,i,d}=k_{t,i,d}h_{d}$). This ranking approach leads to an efficient utilization of the budget and controls the CPA (which is, in fact, a measure of efficiency), but the multiple-choice nature of the problem adds complexity. Algorithm 1 maintains the structure of standard knapsack heuristics (Dantzig, 2002; Kellerer et al., 2004), with two main blocks: advertisement slot ranking by efficiency (lines 3-12) and accumulation of slots until the budget is over (lines 14-24). However, it presents three key modifications:
@@ 4.3. Oracle-slot algorithm
Exposure p. slot 1 $\bm{h_{1}}$ 1 1 1 1
Exposure p. slot 2 $\bm{h_{2}}$ 0.8 0.8 0.8 0.8
Cost slot 1 $\bm{k_{1}}$ 1 1 1 1
Cost slot 2 $\bm{k_{2}}$ 0.36 0.84 0.36 0.84
Conversion p. $\bm{\mu}$ 0.1 0.04 0.2 0.04
Mean cost slot 1 $\bm{\tilde{k}_{1}=k_{1}h_{1}}$ 1 1 1 1
Mean cost slot 2 $\bm{\tilde{k}_{2}=k_{2}h_{2}}$ 0.3 0.7 0.3 0.7
Mean conv. slot 1 $\bm{\tilde{\mu}_{1}=\mu h_{1}}$ 0.1 0.04 0.2 0.04
Mean conv. slot 2 $\bm{\tilde{\mu}_{2}=\mu h_{2}}$ 0.08 0.032 0.16 0.032
Efficiency slot 1 $\bm{\eta_{1}=\mu/k_{1}}$ 0.1 0.04 0.2 0.04
Efficiency slot 2 $\bm{\eta_{2}=\mu/k_{2}}$ 0.278 0.048 0.556 0.048
@@
Eff. upgrade 3 $\rightarrow$ 2 $\bm{\tilde{\eta}_{2}=\eta_{2}}$ 0.278 0.048 0.556 0.048
@@ 4.4. Oracle-upgrade algorithm
The oracle-slot algorithm suffers from one limitation: maybe counter-intuitively, adding the most efficient slot currently not in the solution set $R^{*}$ is not always the most efficient way to extend $R^{*}$. We clarify this statement with an example. For simplicity, let us consider two IOs with two advertisement slots each (D=2). We call $S_{i,j}$ the slot $j$ of IO $i$. The specifications of the two IOs are detailed in Table 1. According to the efficiency values of Example 1, we have that $S_{1,2}$ is the most efficient, with $\eta_{1,2}=0.278$, followed by $S_{1,1}$ ($\eta_{1,1}=0.1$), $S_{2,2}$ ($\eta_{2,2}=0.048$) and $S_{2,1}$ ($\eta_{2,1}=0.4$). And in fact, when bidding proportionally to the conversion probability, we would win slot $S_{1,2}$ with a bid coefficient $\alpha$ such that $1/\eta_{1,2}\leq\alpha<1/\eta_{1,1}$, $S_{1,1}$ with $1/\eta_{1,1}\leq\alpha<1/\eta_{2,2}$, $S_{1,1}$ and $S_{2,2}$ with $1/\eta_{2,2}\leq\alpha<1/\eta_{2,1}$ and finally $S_{1,1}$ and $S_{2,1}$ with $\alpha\geq 1/\eta_{2,1}$. However, we can easily verify that this strategy does not always achieve the maximum expected conversions for a given budget. In fact, winning $S_{1,1}$ leads to $0.1$ expected conversions for an expected cost of $1$, while winning $S_{1,2}$ and $S_{2,2}$ leads to $0.112$ expected conversions, for the same expected cost. This inefficiency is caused by $S_{1,2}$ being much more efficient than $S_{1,1}$. Therefore, although $S_{1,1}$ is more efficient than $S_{2,2}$, it is still better to win $S_{2,2}$ than $S_{1,1}$, in order not to lose the very efficient $S_{1,2}$ (reminder: we cannot win both $S_{1,1}$ and $S_{1,2}$, because they are slots of the same IO).
@@ 6. Experiments and results
We trained auto-bidding agents using OIL in the advertisement campaign simulator built on the advertisement traffic from the Alimama dense and sparse datasets. Budget and target CPA values for each simulated campaign were extracted from uniform distributions encompassing the dataset ranges. To capture campaign dynamics and temporal patterns from past auctions, we engineered 60 features (see Appendix) and fed them into a three-layer fully connected network with 256 units per layer to generate bids. Each OIL training experiment involved 10 million environment interactions, which were sufficient for convergence (Fig. 4). Detailed hyperparameters are listed in the Appendix. We evaluated three versions of OIL:
@@
1. Online LP. A learning-free optimization method, implemented by Xu et al. (Xu et al., [n. d.]) based on ideas of Aggarwal et al. (Aggarwal et al., 2019). This algorithm uses Linear Programming to compile a table of the best bids for different values of budget and target CPA, which is used to decide how much to bid at test time.
2. Behavior cloning (BC). A simple algorithm mimicking bids from the Alimama dataset.
@@
4. Proximal Policy Optimization (PPO). A robust online RL algorithms (Schulman et al., 2017), implemented in Stable Baselines 3 (Raffin et al., 2021), which we evaluated in three versions: as in IQL, we considered two different reward functions (PPO-score and PPO-shaped) and we also allowed the agent to output a two-slope action (PPO-shaped-2s), as we did with OIL-upgrade-2s. This was not possible with IQL, because the bids of the policies featured in the Alimama are linearly proportional to the expected conversion (single slope).
@@
As shown in Table 2, OIL-slot outperformed all baselines, achieving a 21.2% and 17.83% relative score improvement on Alimama-dense and Alimama-sparse datasets, respectively. Notably, OIL-slot balanced high conversion counts and near-complete budget utilization with low CPA. Although IQL occasionally achieved lower CPA than OIL, it consistently sacrificed almost 50% of the conversions, resulting in lower overall score.
To validate our online training choice, we trained OIL-slot offline with the Alimama dataset, replacing agent bids with oracle bids. While the offline-trained agent outperformed all baselines (score: $422.43\pm 6.12$ dense, $31.48\pm 0.18$ sparse), its score was over 10% lower than the online version (score: $466.36\pm 0.43$ dense, $35.29\pm 0.05$ sparse), confirming the advantages of online training.
Figure 4. Learning curves PPO and OIL, Alimama-dense (left) and Alimama-sparse (right).

### v3

@@ 1. Introduction
3. Experimental validation. Using a simulator based on the AuctionNet dataset (Su et al., 2024), we show that OIL-trained agents achieve near-oracle performance and surpass other approaches, including PPO, BC, IQL, and LP.
@@ 2. Related work
Auto-bidding in second price auctions. A core component of advertisement bidding is estimating the utility value $V=K\times\mu$ of each IO, where $K$ is the target CPA and $\mu$ is the acquisition rate (Xu et al., 2016). $K$ is a parameter selected by the advertiser, stating how much it is ready to pay for an acquisition, while $\mu$ can be a function of the advertiser, the potential customer and the context (Yan et al., 2009; Wu et al., 2009; Yuan et al., 2013; Zhang et al., 2014). It can be either estimated by the advertisement platform or by the advertiser itself (Diemert Eustache, Meynet Julien et al., 2017). $V$ defines the fair value of an IO (Xu et al., 2016). Theoretical results prove that bidding equal to the fair value (truth-telling) is optimal for second-price auctions, under the assumption of infinite budget (Krishna, 2009; Roughgarden, 2010). However, when the budget is limited, it might be convenient to bid below the fair value, to increase the acquisition efficiency (Gong et al., 2023), or even above the fair value, if the advertiser can accept exceeding the target CPA to increase the number of acquisitions. The problem is more complex in the presence of multiple advertisement slots for a single IO. In this case, the efficiency of these slots varies with the position, and techniques such as bid shading can be employed to maximize the number of acquisitions (Zhang et al., 2021; Gong et al., 2023). Previous work has focused on optimal advertisement bidding when the fair value of an IO is unknown (Ren et al., 2017; Gligorijevic et al., 2020; Bompaire et al., 2021). In this work, instead, we consider the framework defined by Su et al. (Su et al., 2024), in which the advertisement platform provides an estimation of the acquisition probability, and the advertiser optimizes its bids accordingly, as a function of its budget, target CPA and its understanding of the advertisement traffic patterns. We chose this specific problem because it can be tackled using the recently released AuctionNet dataset, which, with its more than 500 million recorded bids (Su et al., 2024), is the largest open-source advertisement bidding dataset to this date.
@@ 3. Dataset and simulator
Figure 2. Example IOs from the AuctionNet-dense dataset. A. Conversion probabilities for a single advertiser and advertisement campaign. B. Conversion probability at one time step (red dotted line, A) as a function of their lowest price (3rd slot). IOs above the dotted line have an expected CPA lower than the target ($K=8$).
Open-source bidding datasets have historically been limited in size and feature diversity (Ou et al., 2023), and usually focus on predicting the click and conversion probability (Diemert Eustache, Meynet Julien et al., 2017). Consequently, no common benchmark for auto-bidding agents exists, and recent works have relied on private datasets and closed-source algorithm implementations (Wang et al., 2017a; Zhang et al., 2021; Zhou et al., 2021; He et al., 2021; Wang et al., 2022; Wen et al., 2022). This tendency complicates the comparison between methods (Ou et al., 2023). However, recently Alimama has released a large scale dataset with more than 500 million advertisement bids showcasing realistic advertisement traffic patterns (Su et al., 2024).
Dataset. The dataset comprises two datasets, AuctionNet-dense and AuctionNet-sparse, spanning 21 campaign periods of 48 time steps each. For each time step and IO, bids from 48 advertisers are recorded, competing for three advertisement slots assigned via second-price auction. This means that the advertisers presenting the three highest bids win the three slots, paying a price equal to the next highest bid (second to fourth). The main distinction between the two datasets is the average conversion probability, which is one order of magnitude lower in the sparse dataset, resulting in fewer conversions (hence the name ”sparse”). Each advertiser is characterized by a fixed total campaign budget and a fixed target CPA, and bids according to its specific strategy. The dataset also provides the expected conversion probability for each IO, that was made available to the advertisers by the platform at bidding time. Conversion probabilities differ across advertisers for the same IO, reflecting varying product appeal to different potential customers. Advertisers face stochastic advertisement traffic (Fig. 2 A). To avoid both overbidding (and missing out on valuable IOs later due to lack of budget) and underbidding (and ending up with unspent budget that they had reserved for valuable IOs which never materialized), the advertisers need to adjust their strategy to the advertisement traffic they expect in the future. The conversion probabilities of the IOs at the same time step and for the same advertiser present a fairly spread distribution (Fig. 2 B), with only a fraction being convenient for the advertiser (i.e., the bid necessary to win the IO leads to a lower CPA than the target).
@@
Simulator. To train auto-bidding agents, we developed an advertisement campaign simulator based on the AuctionNet dataset (Fig. 3). The process is as follows: 1) the auto-bidding agent impersonates one of the 48 advertiser, sampled at random, and is assigned a budget and a target CPA. 2) At each time step, the simulator provides estimated conversion probabilities for available IOs (typically 1000 10,000 per time step). 3) The auto-bidding agent places bids without observing those of the competitors (closed-envelope auction). 4) Competitor bids are extracted from the dataset, assuming that they are not influenced by the auto-bidding agent’s behavior. 5) Slot allocation and pricing are determined via second-price auction.
@@ 4.2. Heuristic algorithm
Exact solutions to Problem 2 are computationally infeasible for campaigns involving hundreds of thousands of IOs, such as the ones in the AuctionNet dataset. Instead, we employ a greedy heuristic approximation. Inspired by knapsack problem solutions, the algorithm ranks slots by efficiency (e.g., conversion probability over cost, $\mu_{t,i}/k_{t,i,d}$) and selects slots until the budget is exhausted. Expected conversion probabilities and costs are adjusted for exposure probabilities since conversions and costs only occur when ads are displayed (effective conversion probability $\tilde{\mu}_{t,i,d}=\mu_{t,i}h_{d}$ and effective cost $\tilde{k}_{t,i,d}=k_{t,i,d}h_{d}$). This ranking approach leads to an efficient utilization of the budget and controls the CPA (which is, in fact, a measure of efficiency), but the multiple-choice nature of the problem adds complexity. Algorithm 1 maintains the structure of standard knapsack heuristics (Dantzig, 2002; Kellerer et al., 2004), with two main blocks: advertisement slot ranking by efficiency (lines 3-12) and accumulation of slots until the budget is over (lines 14-24). However, it presents three key modifications:
@@ 4.3. Oracle-slot algorithm
Exposure p. slot 1 $\bm{h_{1}}$ 1.000 1.000 1.000 1.000
Exposure p. slot 2 $\bm{h_{2}}$ 0.800 0.800 0.800 0.800
Cost slot 1 $\bm{k_{1}}$ 1.000 1.000 1.000 1.000
Cost slot 2 $\bm{k_{2}}$ 0.375 0.875 0.375 0.875
Conversion p. $\bm{\mu}$ 0.100 0.040 0.200 0.040
Mean cost slot 1 $\bm{\tilde{k}_{1}=k_{1}h_{1}}$ 1.000 1.000 1.000 1.000
Mean cost slot 2 $\bm{\tilde{k}_{2}=k_{2}h_{2}}$ 0.300 0.700 0.300 0.700
Mean conv. slot 1 $\bm{\tilde{\mu}_{1}=\mu h_{1}}$ 0.100 0.040 0.200 0.040
Mean conv. slot 2 $\bm{\tilde{\mu}_{2}=\mu h_{2}}$ 0.080 0.032 0.160 0.032
Efficiency slot 1 $\bm{\eta_{1}=\mu/k_{1}}$ 0.100 0.040 0.20 0.040
Efficiency slot 2 $\bm{\eta_{2}=\mu/k_{2}}$ 0.267 0.046 0.533 0.046
@@
Eff. upgrade 3 $\rightarrow$ 2 $\bm{\tilde{\eta}_{2}=\eta_{2}}$ 0.267 0.046 0.533 0.046
@@ 4.4. Oracle-upgrade algorithm
The oracle-slot algorithm suffers from one limitation: maybe counter-intuitively, adding the most efficient slot currently not in the solution set $R^{*}$ is not always the most efficient way to extend $R^{*}$. We clarify this statement with an example. For simplicity, let us consider two IOs with two advertisement slots each (D=2). We call $S_{i,j}$ the slot $j$ of IO $i$. The specifications of the two IOs are detailed in Table 1. According to the efficiency values of Example 1, we have that $S_{1,2}$ is the most efficient, with $\eta_{1,2}=0.278$, followed by $S_{1,1}$ ($\eta_{1,1}=0.100$), $S_{2,2}$ ($\eta_{2,2}=0.048$) and $S_{2,1}$ ($\eta_{2,1}=0.040$). And in fact, when bidding proportionally to the conversion probability, we would win slot $S_{1,2}$ with a bid coefficient $\alpha$ such that $1/\eta_{1,2}\leq\alpha<1/\eta_{1,1}$, $S_{1,1}$ with $1/\eta_{1,1}\leq\alpha<1/\eta_{2,2}$, $S_{1,1}$ and $S_{2,2}$ with $1/\eta_{2,2}\leq\alpha<1/\eta_{2,1}$ and finally $S_{1,1}$ and $S_{2,1}$ with $\alpha\geq 1/\eta_{2,1}$. However, we can easily verify that this strategy does not always achieve the maximum expected conversions for a given budget. In fact, winning $S_{1,1}$ leads to $0.1$ expected conversions for an expected cost of $1$, while winning $S_{1,2}$ and $S_{2,2}$ leads to $0.112$ expected conversions, for the same expected cost. This inefficiency is caused by $S_{1,2}$ being much more efficient than $S_{1,1}$. Therefore, although $S_{1,1}$ is more efficient than $S_{2,2}$, it is still better to win $S_{2,2}$ than $S_{1,1}$, in order not to lose the very efficient $S_{1,2}$ (reminder: we cannot win both $S_{1,1}$ and $S_{1,2}$, because they are slots of the same IO).
@@ 6. Experiments and results
We trained auto-bidding agents using OIL in the advertisement campaign simulator built on the advertisement traffic from the AuctionNet dense and sparse datasets. Budget and target CPA values for each simulated campaign were extracted from uniform distributions encompassing the dataset ranges. To capture campaign dynamics and temporal patterns from past auctions, we engineered 60 features (see Appendix) and fed them into a three-layer fully connected network with 256 units per layer to generate bids. Each OIL training experiment involved 10 million environment interactions, which were sufficient for convergence (Fig. 4). Detailed hyperparameters are listed in the Appendix. We evaluated three versions of OIL:
@@
1. Online LP. A learning-free optimization method, implemented by Su et al. (Su et al., 2024) based on ideas of Aggarwal et al. (Aggarwal et al., 2019). This algorithm uses Linear Programming to compile a table of the best bids for different values of budget and target CPA, which is used to decide how much to bid at test time.
2. Behavior cloning (BC). A simple algorithm mimicking bids from the AuctionNet dataset.
@@
4. Proximal Policy Optimization (PPO). A robust online RL algorithms (Schulman et al., 2017), implemented in Stable Baselines 3 (Raffin et al., 2021), which we evaluated in three versions: as in IQL, we considered two different reward functions (PPO-score and PPO-shaped) and we also allowed the agent to output a two-slope action (PPO-shaped-2s), as we did with OIL-upgrade-2s. This was not possible with IQL, because the bids of the policies featured in the AuctionNet are linearly proportional to the expected conversion (single slope).
@@
As shown in Table 2, OIL-slot outperformed all baselines, achieving a 21.2% and 17.83% relative score improvement on AuctionNet-dense and AuctionNet-sparse datasets, respectively. Notably, OIL-slot balanced high conversion counts and near-complete budget utilization with low CPA. Although IQL occasionally achieved lower CPA than OIL, it consistently sacrificed almost 50% of the conversions, resulting in lower overall score.
To validate our online training choice, we trained OIL-slot offline with the AuctionNet dataset, replacing agent bids with oracle bids. While the offline-trained agent outperformed all baselines (score: $422.43\pm 6.12$ dense, $31.48\pm 0.18$ sparse), its score was over 10% lower than the online version (score: $466.36\pm 0.43$ dense, $35.29\pm 0.05$ sparse), confirming the advantages of online training.
Figure 4. Learning curves PPO and OIL, AuctionNet-dense (left) and AuctionNet-sparse (right).